OpenStudy (anonymous):

Find the solution to the system of equations by using either graphing or substitution. y=4x and y+x=5 a. (-4,1) b (1,4) c. (-3,2) d. (2,3)

jimthompson5910 (jim_thompson5910):

y+x=5 4x+x=5 ... replace y with 4x (since y = 4x)

jimthompson5910 (jim_thompson5910):

Are you able to solve 4x+x=5 for x?

OpenStudy (anonymous):

B. https://www.desmos.com/calculator/mxk5mhwgea